Two particles P and Q are initially 40 m apart P behind Q. Particle P starts moving with a uniform velocity 10 m/s towards Q. Particle Q starting from rest has an acceleration 2 m/s² in the direction of velocity of P. Then the minimum distance between P and Q will be?

Grade:11

2 Answers

Vikas TU
7 years ago
For Partcle P,
let it covered the distance x then,
x + 40 = 10*t.............(1)
 
For Particle Q,
x  = 0 + 0.5*2*t^2................(2)
solve the both eqns. and get t and x values.
Mohit
6 years ago
The minimum distance will be found when Q moves with 10m/s velocity which can be found byv=u+at10=0+2tt=5 sec.Distance traveled by P in 5 seconds is 5*10 i.e 50m. And distance traveled by Q in 5 seconds iss=0.5*2*(5)^2s=25m.So minimum distance between P and Q will be 40+25-50=15m